OpenStudy (anonymous):

Evaluate polynomial for x = −3. 4x2 − 6x + 9 for x = -3 4(-3)^2 – 6(-3) + 9 = 4(-9) – 6(-3) + 9 = 36+ 18 + 9 = -3 Can someone check my problem for me please

OpenStudy (anonymous):

when u square a number u square the negative sign as well -3 ur step 2 is wrong and 63 should be answer

OpenStudy (anonymous):

= 4(-9) – 6(-3) + 9 so what is wrong in this step?

OpenStudy (anonymous):

(-3)^2=9

OpenStudy (anonymous):

ok mathematicalro..where to from there. I'm confused

OpenStudy (anonymous):

Would any one mind writing this out the correct way so I can understand?

OpenStudy (anonymous):

its not 4(-9) – 6(-3) + 9 its 4(9)-6(-3)+9 = 36+ 18 + 9 =63